OpenStudy (anonymous):

simplify: y^2-7y-18/y^2-4y-12 divided by y^2+8y+15/y^2-y-30

OpenStudy (anonymous):

im so aggravated that im getting hung up on these

OpenStudy (callisto):

y^2-7y-18/y^2-4y-12 = (y+2)(y-9) / (y+2)(y-6) = (y-9)/(y-6) y^2+8y+15/y^2-y-30 = (y+3)(y+5) / (y+5)(y-6) = (y+3)/(y-6) therefore y^2-7y-18/y^2-4y-12 divided by y^2+8y+15/y^2-y-30 = [(y-9)/(y-6)] / [ (y+3)/(y-6)] = (y-9)/(y+3)